OpenStudy (anonymous):

Discuss the continuity of the function f, where f is defined by f(x) = 2x, if x < 0 0, if 0<=x<=1 4x, if x>1

OpenStudy (calculusfunctions):

A function f(x) is continuous at a point x = a iff i). f(a) is defined, ii). limit of f(x) exists as x "approaches" a and equals L This is true iff\[\lim_{x \rightarrow a ^{-}}f(x)=\lim_{x \rightarrow a ^{+}}f(x)=L\]iii).\[\lim_{x \rightarrow a}f(x)=f(a)\]If any of these conditions are not met, then f(x) is discontinuous at x = a. There are three types of discontinuities: I). Removable Discontinuity: i). f(a) may or may not be defined, ii). limit of f(x) exists as x "approaches" a and equals L, whether f(a) is defined or not. However if f(a) is defined, then limit of f(x) as x "approaches a equals L. II). Infinite Discontinuity: i). f(a) is NOT defined, ii). limit of f(x) as x "approaches" a from the left or the right equals negative or positive infinity. In other words, the limit does not exist as x "approaches" a. Note: vertical asymptotes are infinite discontinuities. III). Jump Discontinuity: i). f(a) may or may not be defined, ii). limit of f(x) as x "approaches" DOES NOT EXIST. However for this type of discontinuity, there arises a special case. If f(a) is defined, then iff i).\[\lim_{x \rightarrow a ^{-}}f(x)=f(a)\]then the function f(x) is continuous from the left, or if ii). \[\lim_{x \rightarrow a ^{+}}f(x)=f(a)\]then f(x) is continuous from the right.

OpenStudy (calculusfunctions):

CORRECTION: PLEASE NOTE THE FOLLOWING ERROR!!! FOR REMOVABLE DISCONTINUITIES, f(a) is NOT DEFINED but he LIMIT DOES EXIST!!!!